How many gallons of a 12%-salt solution must be mixed with 6 gallons of a 21%-salt solution to obtain a 15%-salt solution?
Answer by Maths68(1474)

Solution A
Amount = x gallons
Concentration =12% =0.12

Solution B
Amount = 6 gallons
Concentration =21% = 0.21

Resultant Solution
Amount =6+x gallons
Concentration =15%=0.15

[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant


(x)(0.12) + (6)(0.21) = (6+x)(0.15)
0.12x+1.26=0.9+0.15x
0.12x-0.15x=0.9-1.26
-0.03x=-0.36
-0.03x/-0.03=-0.36/0.03
x=12



12 gallons of a 12%-salt solution must be mixed with 6 gallons of a 21%-salt solution to obtain 18 gallons of a 15%-salt solution?